Home‎ > ‎Crew 164‎ > ‎


> SOL: 11Person filling out
Summary Title: Let Us Stay
> On Mars!
> Mission Status: The overall operations mission
> performance is nominal and routine operations are
> on-going.
> Sol Activity Summary: EVA#11 A and B.
> Disassembling the CRV. End of mission report
> writing.
> Look Ahead Plan: Last EVA, end of mission report,
> getting ready to host 12 people on Saturday (next crew and
> film school students)
> Anomalies in work: At the time of writing this
> report, the observatory shutter was not yet tested after the
> day of battery charging. The solar panel was pointing South
> all day.Weather: Broken clouds. Warm day. Dry
> ground. Low wind.
> Crew Physical Status: No issue.
> EVA: EVA in the morning to test the CRV in canyon
> conditions. In the afternoon the 3 geologists did the RMR
> Geology experiment again.
> Reports to be filed: - Commander
> report- Engineering report- Journalist
> report - EVA #11A report- EVA #11B
> report- EVA #12 request
> Support Requested: /-- 
> Mohammad Iranmanesh